RC meeting and Run Plan, Tuesday November 8th

From SBS wiki
Jump to: navigation, search

Return to the RC Meeting Calendar

RC Meeting Information

Time:at 3:30pm today

  • Location: Counting house conference room

Useful Links

Current Status

Hall Status

  • Target Cell "Windmill" was installed overnight by the target team.
  • Hall in restricted and Jessie's team finished hall checkout today. BigBite is at 36.5 degrees.
  • We are currently waiting for sweep of the hall and transition to beam permit.
  • Possible updates from Andrew.

Open Issues

  • BB angle at 36.5 degrees matching print. Is this acceptable?
  • Moller HV is working. Thanks to Mark Taylor and Don for the quick fix.
  • Simona replaced the raster scope.
  • Nitrogen for cooling jets were adjusted today by Todd and final target adjustments and checks complete.
  • Reference cell gas for ion chamber calibrations. Open discussion.
  • Beam restoration ops tonight. Update from Simona.
  • Oven readback appears to be inaccurate. Update from Todd.

Polarized Target

  • Current Polarization: N/A (Will begin polarization after sweep of the hall)
  • Hourly screenshots suspended until be begin polarizing the target. Lasers locked out.
  • Possible update from Arun.

Subsystems

  • Any Updates?

Future Plans

  • Hall currently in restricted access until sweep is complete.
  • Following sweep, we will go to beam permit and begin beam restoration process led by Simona.
  • Target cell polarization will begin during swing.

Other Topics

Experimental progress (for MCC planning meeting)

  • 6.1% of accumulated charge for PRODUCTION data (13.55 Coulombs out of 232)
  • 31% Experimental run time (44 out of 141 days)

Run Plan After Configuration Change

Continue to update 'standard' run plan for 3-pass data taking as necessary.

Current Plan:

  • GEM luminosity studies could be performed between beam restoration and target polarization. See procedure.
  • NMR measurement
  • Restore beam
  • Optics and H2 Ref Cell (Sometime during run)
  • Production Run Plan
    • 1.) He-3 runs, at 45uA of beam current. Use ps3==0, ps5=12, all others -1. Take four runs that should each last for one hour.
    • 2.) Move the target to the pickup coil position, and the TO should run an NMR sweep to measure the polarization.
    • 3.) Return to step 1.)

While on shift:

  • Keep an eye out for drifting BPM positions
  • Log all DAQ crashes in the excel worksheet along with the reason for the crash
  • If beam has been down due to accelerator issues perform a checkout of the beam profile by doing harp scans and check beam centering on target using the Carbon hole target, see Simona's procedure: procedure, If in doubt about whether you should go through the beam checkout procedure, call this number: 7578693131.
  • Cross check the rate with what is expected: log
  • Cross check with the expected ion chamber readings: log

Notices in Effect for Shift Crew

All SL shifts are on standby today

Be sure to follow these procedures:

Shifts and Long Term Issues

  • Shift Openings: We're looking good for now! Schedule

Planned Accesses to the Hall

In restricted Access. Survey and alignment ongoing.

Special Requests

Nothing scheduled at present.